# Runtime Analysis of the SMS-EMOA for Many-Objective Optimization

### 16 Dec 2023 v1

@@ 6.2 Runtime
Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
With similar proof ideas in Lemma 5 to 7 but noticing the probability of generating a solution with a specific Hamming distance, we obtain the following lemmas for reaching the first inner Pareto optimum (Lemma 17), covering all inner Pareto front points (Lemma 18), and covering the remaining Pareto front points (Lemma 19).
Lemma 17.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $C_{n/2}^{\beta}\binom{n}{k}mk^{\beta}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
Proof.
We use similar arguments of Lemma 5 except following probability calculations. The probability to generate a $y$ with $|y_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ from $x$ with $A_{i}<k$ or with $A_{i}>n^{\prime}-k$ is at least $\binom{n}{k}^{-1}P_{k}^{\beta}\geq\binom{n}{k}^{-1}\mathopen{}\mathclose{{}\left(C_{n/2}^{\beta}}\right)^{-1}k^{-\beta}$. Hence such $y$ is generated in expected $\binom{n}{k}C_{n/2}^{\beta}k^{\beta}$ iterations. Hence the expected iterations to reach an inner Pareto optimum is at most $\frac{m}{2}\binom{n}{k}C_{n/2}^{\beta}k^{\beta}=\binom{n}{k}C_{n/2}^{\beta}mk^{\beta}$ iterations. ∎
Lemma 18.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population contains at least one inner Pareto optimum. Then after at most $\frac{e\beta}{\beta-1}n\mu M$ iterations in expectation, all inner Pareto front points are covered.
Proof.
We use similar arguments of Lemma 6 except following probability calculation. The probability to generate an offspring with its function value of an uncovered inner Pareto front point from a neighbor covered by the population is at least $\frac{1}{n\mu}P_{1}^{\beta}=\frac{\beta-1}{e\beta n\mu}$. ∎
Lemma 19.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $\binom{n}{k}C_{n/2}^{\beta}\mu Mk^{\beta}$ iterations in expectation, the full Pareto front is covered.
Proof.
We use similar arguments of Lemma 7 except following probability calculation. The probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
$\frac{1}{\mu}\binom{n}{k}^{-1}P_{k}^{\beta}\geq\binom{n}{k}^{-1}\mathopen{}\mathclose{{}\left(C_{n/2}^{\beta}}\right)^{-1}\frac{1}{\mu}k^{-\beta}.\qed$
Considering the whole process for the full coverage of the Pareto front from Lemmas 17 to 19, we have the following runtime result for the SMS-EMOA with heavy-tailed mutation.
@@
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $O(M\mu k^{\beta-0.5-k}n^{k})$ iterations or function evaluations in expectation, the full Pareto front is covered.
Theorem 20 shows the best asymptotic upper bound of the runtime of the SMS-EMOA with heavy-tailed mutation is $O(M^{2}n^{k}k^{\beta-0.5-k})$. Comparing with the runtime of $O(M^{2}n^{k})$ for $\mu=\Theta(M)$ for the original SMS-EMOA from Theorem 8, we see a speed-up of $k^{k+0.5-\beta}$, which is also the speed-up reached for the single objective and two objectives.
7 SMS-EMOA Also Performs Well for Bi-objective Optimization
The above sections discussed the runtime of the SMS-EMOA for many objectives, and showed that it performs well for many objectives, which is different from the original NSGA-II. Since the only theory paper [BZLQ23] merely considers its performance for the bi-objective OJZJ, in this section, we will analyze its runtime for other well-known bi-objective benchmarks to have a more comprehensive theoretical understanding.
7.1 OneMinMax and LOTZ
OneMinMax [GL10] (also the CountingOnesCountingZeroes [${}^{+}$02]) and LOTZ [LTZ04] (also the weighted version WLPTNO [QYZ13]) are the bi-objective counterpart of the well-analyzed single-objective OneMax and LeadingOnes respectively. For OneMinMax, two objectives count the number of ones and zeros in a given bit-string respectively. For LOTZ, the first objective counts the continuous ones starting from the first bit position, and the second one counts the continuous zeros starting from the last bit position. Both benchmarks with a simple and clear structure facilitate the fundamental theoretical understanding of the MOEAs. We note that there are other benchmarks [QTZ16, LZZZ16, DOSS23b], but here we focus on the two kinds of most popular benchmarks for analysis.
Here are formal definitions of the OneMinMax and LOTZ.
Definition 21 ([GL10]).
@@ 7.1 OneMinMax and LOTZ
$\textsc{OneMinMax}(x)=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}(1-x_{i}),\sum_{i=1}^{n}x_{i}}\right)$
@@
$\textsc{LOTZ}(x)=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}\prod_{j=1}^{i}x_{j},\sum_{i=n}^{1}\prod_{j=n}^{i}x_{j}}\right)$
@@
7.2 Runtime
We now establish the runtime of the SMS-EMOA on the OneMinMax and LOTZ. Note that the Pareto front sizes for OneMinMax and LOTZ are $n+1$. Hence, with the similar proof of Lemma 4, we know that for $\mu\geq n+1$ and for a specific function value $u$ reached by $F_{1}$, one individual in the non-empty $\cup_{i\geq 2}F_{i}$ or one individual $x$ in $F_{1}$ with $\Delta(x,F_{1})=0$ will be removed and $u$ will remain covered in the next generation. We have the following formal description.
Lemma 23.
Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ or OneMinMax problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
Theorem 24 shows the runtime of SMS-EMOA on OneMinMax. The proof considers the probability to reach an uncovered Pareto front point from a specific covered one. From Lemma 23 that any covered Pareto front point will be kept, we then turn such probability to the runtime argument.
@@ 7.2 Runtime
Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the OneMinMax problem with problem size $n$. Then after at most $2e\mu n(\ln n+1)$ iterations in expectation, the full Pareto front is covered.
Proof.
Since each search point in $\{0,1\}^{n}$ is Pareto optimal, from Lemma 23, we know that any Pareto front point covered by the population $P_{t}$ will be maintained in all future generations. Assume that $(i,n-i)$ for some $i\in[0..n]$ is covered in $P_{t}$. Then the probability to reach a solution with function value $(i+1,n-i-1)$ is at least
$\frac{1}{\mu}\frac{n-i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{n-i}{e\mu n},$
and the probability to reach a solution with function value $(i-1,n-i+1)$ is at least
$\frac{1}{\mu}\frac{i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{i}{e\mu n}.$
Thus the expected numbers of iterations to cover $(i+1,n-i-1)$ and $(i-1,n-i+1)$ are at most $\frac{e\mu n}{n-i}$ and at most $\frac{e\mu n}{i}$ respectively. Hence, the expected number of iterations to cover the full Pareto front is
$\Bigg{(}\sum_{j=i}^{1}\frac{e\mu n}{n-j}\Bigg{)}+\mathopen{}\mathclose{{}\left(\sum_{j=i}^{n-1}\frac{e\mu n}{j}}\right)\leq 2\sum_{j=1}^{n-1}\frac{e\mu n}{j}\leq 2e\mu n(\ln n+1).\qed$
We now establish the runtime of SMS-EMOA on LOTZ, see Theorem 25. In the proof, the process is divided into two stages. The first stage is to cover a Pareto front point for the first time and the second stage is to cover the full Pareto front starting from one covered Pareto front point. For the first stage, we pessimistically consider the time to reach $1^{n}$ and for the second stage, we consider the time to reach $01^{i}0^{n-i},i=n-1,\dots,$ 0 one by one.
@@
Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ problem with problem size $n$. Then after at most $2e\mu n^{2}$ iterations in expectation, the full Pareto front is covered.
Proof.
We first consider the time to reach $1^{n}$ in the first stage. Let $S=\arg\max_{x\in P_{t}}f_{1}(x)$ and $S^{\prime}=\arg\max_{x\in R_{t}}f_{1}(x)$. Then there is at least one individual $x\in S^{\prime}$ such that $x\in F_{1}$ (If not, there is $y\notin S^{\prime}$, such that $x\prec y$ for all $x\in S^{\prime}$. That is, $f_{1}(y)\geq f_{1}(x),f_{2}(y)\geq f_{2}(x),$ and at least one of the inequality is strict. Since $y\notin S$, we know that $f_{1}(y)<f_{1}(x)$, which contradicts $f_{1}(y)\geq f_{1}(x)$). From Lemma 23, we know that $f(x)$ (also $f_{1}(x)$) will be covered by $P_{t+1}$. Hence, to reach $1^{n}$, we only need to reach such $S^{\prime}$ with $f_{1}(S^{\prime})=n$. We calculate
$\Pr[f_{1}(S^{\prime})>f_{1}(S)\mid f_{1}(S)]\geq\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}$
where we consider choosing one individual of $S$ and flipping the critical bit value from 00 to $1$. Since $f_{1}(S)$ will not decrease, we know that the expected iterations to witness such $S^{\prime}$ is at most $e\mu n$. Since there are at most $n$ such $S^{\prime}$ with $f_{1}(S^{\prime})>f_{1}(S)$ to reach $1^{n}$, we know that $1^{n}$ will be reached in at most $e\mu n^{2}$ in expectation.
Once $1^{n}$ is reached, we consider the time to reach $1^{n-1}0,\dots,10^{n-1},0^{n}$ one by one. Assume that $1^{i}0^{n-i}$ with $i\in[1..n]$ is reached. Then to reach $1^{i-1}0^{n-i+1}$, we only need to choose $1^{i}0^{n-i}$ from $P_{t}$, and flip the $i$-th bit from $1$ to 00 with other bits unchanged, which happens with probability
$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}.$
From Lemma 23, we know that any Pareto front point will be maintained in all future generations once reached. Hence, $1^{i}0^{n-i}\in P_{\tau}$ for all $\tau\geq t$. Thus, the expected number of iterations to reach $1^{i-1}0^{n-i+1}$ is at most $e\mu n$. Since we have $n$ other Pareto optima to reach, we know that the expected number of iterations to cover all remaining Pareto front points is at most $e\mu n^{2}$.
In summary, the full Pareto front will be covered in at most $2e\mu n^{2}$ iterations in expectation. ∎
Note that the runtime of the GSEMO and NSGA-II are $O(n^{2}\ln n)$ and $O(n^{3})$ for the OneMinMax and LOTZ respectively. Hence, Theorems 24 and 25 shows the SMS-EMOA achieves the same asymptotic runtime as GSEMO and NSGA-II.
8 Conclusion
Motivated by the observed difficulty of the NSGA-II for many objectives, this paper resorted to the SMS-EMOA, a variant of the steady-state NSGA-II, and proved that, different from the NSGA-II, it efficiently solves the $m$OJZJ problem. Noting that the SMS-EMOA also employs non-dominated sorting, but replaces the crowding distance with the hypervolume, this result together with the one of Wietheger and Doerr [WD23] supports the conclusion that non-dominated sorting is a good building block for MOEAs, but the crowding distance has deficiencies for more than two objectives.
We also showed that the stochastic population update proposed in [BZLQ23] for the bi-objective SMS-EMOA becomes less effective for many objectives. All these results in a very rigorous manner support the general knowledge that multi-objective optimization becomes increasingly harder with growing numbers of objectives.

### 25 Jan 2024 v2

@@ Introduction
The assertion of Lemma 4 extends to the SMS-EMOA with heavy-tailed mutation.
With proof ideas similar to those in Lemma 5 to 7, but noticing the different probabilities of generating solutions with a specific Hamming distance, we obtain the following runtime guarantee.
@@
Let $k\leq n^{\prime}/2$ and $\beta>1$. Consider using SMS-EMOA with $\mu\geq M$ and with heavy-tailed mutation to optimize the $m$OJZJ problem. Then after at most $O(M\mu k^{\beta-0.5-k}n^{k})$ iterations or function evaluations in expectation, the full Pareto front is covered.
Compared to the runtime guarantee of Theorem 8 for the original SMS-EMOA, the guarantee of Theorem 17 above for the SMS-EMOA with heavy-tailed mutation is by a factor of asymptotically $k^{k+0.5-\beta}$ stronger.
The SMS-EMOA Also Performs Well for Bi-objective Optimization
The above sections discussed the runtime of the SMS-EMOA for many objectives, and showed that it performs well for many objectives, which is different from the original NSGA-II. Since the only theory paper (Bian et al. 2023) on the SMS-EMOA merely considers its performance on the bi-objective OJZJ problem, to broaden our understanding of this algorithm we now analyze its runtime for the two most prominent bi-objective benchmarks OneMinMax and LOTZ.
OneMinMax and LOTZ
The OneMinMax benchmark introduced by Giel and Lehre (2010) (and also the CountingOnesCountingZeroes benchmark proposed by Laumanns et al. (2002)) are bi-objective counterparts of the famous single-objective OneMax benchmark. Likewise, the LOTZ benchmark defined by Laumanns, Thiele, and Zitzler (2004) (also the weighted version WLPTNO of Qian, Yu, and Zhou (2013)) are bi-objective counterparts of the classic single-objective LeadingOnes benchmark.
For OneMinMax, the two objectives count the number of ones and zeros in a given bit-string, respectively. For LOTZ, the first objective is the number of contiguous ones starting from the first bit position, and the second objective is the number of contiguous zeros starting from the last bit position. Both benchmarks with their simple and clear structure facilitate the fundamental theoretical understanding of MOEAs. We note that there are other benchmarks used in the mathematical analysis of MOEAs (see, e.g. Horoba and Neumann 2008; Brockhoff et al. 2009; Qian, Tang, and Zhou 2016; Li et al. 2016; Dang et al. 2023b), but clearly OneMinMax and LOTZ are the most prominent ones.
Here are the formal definitions of OneMinMax and LOTZ.
Definition 18 (Giel and Lehre (2010)).
@@
$\textsc{OneMinMax}(x)=\left(\sum_{i=1}^{n}(1-x_{i}),\sum_{i=1}^{n}x_{i}\right)$
@@
$\textsc{LOTZ}(x)=\left(\sum_{i=1}^{n}\prod_{j=1}^{i}x_{j},\sum_{i=1}^{n}\prod_{j=i}^{n}(1-x_{j})\right)$
@@
Runtime
We now analyze the runtime of the SMS-EMOA on the OneMinMax and LOTZ benchmarks. Note that the Pareto front, and more generally the size of any set of pairwise non-dominated solutions, for OneMinMax and LOTZ have sizes at most $n+1$. Hence we can use Lemma 4 with $M=n+1$.
Theorem 20 below gives an upper bound for the runtime of SMS-EMOA on OneMinMax. This bound (in terms of fitness evaluations) is of the same asymptotic order as the runtime guarantee for the NSGA-II (Zheng, Liu, and Doerr 2022). Knowing from Lemma 4 that we cannot lose Pareto points, the proof of our results consists of adding the waiting times for finding a new Pareto point from an already existing neighboring one.
@@
Consider using the SMS-EMOA with $\mu\geq n+1$ to optimize the OneMinMax problem with problem size $n$. Then after at most $2e\mu n(\ln n+1)$ iterations (or $\mu+2e\mu n(\ln n+1)$ fitness evaluations) in expectation, the full Pareto front is covered.
Our runtime guarantee for LOTZ, see Theorem 21 below, again coincides with the guarantee for the NSGA-II (Zheng, Liu, and Doerr 2022). In the proof, the process is divided into two stages. The first stage is to cover a Pareto front point for the first time and the second stage is to cover the full Pareto front starting from one covered Pareto front point. For the first stage, we pessimistically consider the time to reach $1^{n}$. For the second stage, we consider the time to generate $01^{i}0^{n-i},i=n-1,\dots,$ 0 one after the other. In both stages, we heavily rely on Lemma 4 asserting that important progress is not lost, which allows to add waiting times of successive progress events.
@@
Consider using the SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ problem with problem size $n$. Then after at most $2e\mu n^{2}$ iterations (or fitness evaluations) in expectation, the full Pareto front is covered.
Note that the runtimes of the GSEMO on OneMinMax and LOTZ are $O(n^{2}\ln n)$ and $O(n^{3})$, respectively (Giel and Lehre 2010; Laumanns, Thiele, and Zitzler 2004). Hence, they agree with our bounds for the SMS-EMOA when using the most interesting population size $\mu=\Theta(n)$.
Conclusion
Motivated by the observed difficulty of the NSGA-II for many objectives, this paper resorted to the SMS-EMOA, a variant of the steady-state NSGA-II, and proved that, different from the NSGA-II, it efficiently solves the $m$OJZJ problem. Noting that the SMS-EMOA also employs non-dominated sorting, but replaces the crowding distance with the hypervolume, this result together with the one of Wietheger and Doerr (2023) supports the conclusion that non-dominated sorting is a good building block for MOEAs, but the crowding distance has deficiencies for more than two objectives.
We also showed that the stochastic population update proposed in (Bian et al. 2023) for the bi-objective SMS-EMOA becomes less effective for many objectives. All these results in a very rigorous manner support the general knowledge that multi-objective optimization becomes increasingly harder with growing numbers of objectives.